# inGRASS: Incremental Graph Spectral Sparsification via Low-Resistance-Diameter Decomposition

### v1

@@ II-B Spectral Graph Decomposition
Furthermore, a similar spectral graph sparsification method is proposed in [9], which decomposes the graph into multiple sets of disjoint cycles using a multilevel spectral graph coarsening framework. By constructing a hierarchy of contracted graphs, the method identifies and adds spectrally-critical edges to the initial sparse graph.
@@ II-C Effective Resistance
$R_{eff}(p,q)=\sum\limits_{i=2}^{N}\frac{w_{p,q}(u_{i}^{\top}b_{pq})^{2}}{\lambda_{i}},$ (2)
@@ III-A Overview of inGRASS Algorithm
This is a one-time operation that equips the inGRASS algorithm to iteratively update the initial sparse graph $H$ upon receiving streams of newly introduced edges in different iterations. In this phase, each node in the initial sparse graph $H^{(0)}$ is assigned a $\log N$-dimensional embedding vector by leveraging low-resistance-diameter (LRD) decomposition. A sparse data structure is then created that provides an efficient estimate of the resistance distance among different pair of nodes in the initial sparse graph $H^{(0)}$.
@@ III-B1 Scalable Estimation of Effective Resistances
Figure 2: A 4-level resistance embedding of the initial sparse graph achieved through the proposed LRD decomposition. Since the embedding vectors for nodes $5$ and $9$ are $[3,3,3,2]^{\top}$ and $[6,2,2,2]^{\top}$, the effective-resistance distance between them is bounded by the resistance diameter of cluster $2$ shown in (d).
@@
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{w_{p,q}(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L_{G}\tilde{u}_{i}},$ (3)
@@ III-B2 Multilevel Resistance Embedding
We utilize a multilevel low-resistance-diameter (LRD) decomposition framework for embedding nodes of the initial sparse graph $H^{(0)}$ into an $O(\log N)$-dimensional space [17]. The embedding process iteratively follows these steps: (S1) efficiently estimate the effective resistance of edges using (3); (S2) contract edges starting from low-resistance-diameter clusters (cluster diameters are initialized to 0 for all nodes in the first level); (S3) replace a contracted edge with a supernode and adjust the resistance diameter of the clusters accordingly.
@@ III-B3 Multilevel Sparse Data Structure
Leveraging the $O(\log N)$-dimensional node embedding vectors obtained in the previous step, we construct a specialized sparse data structure for efficient access to the node clustering indices at different levels. This facilitates efficient storage and retrieval of cluster connectivity information, enabling the identification of spectrally-critical and spectrally-unique edges among the newly added edges. The sparse data structure is promptly updated upon the addition of a newly introduced edge to the sparse graph.
Figure 2 illustrates the computation of node embedding vectors through multilevel LRD decomposition of the initial sparse graph. Increasing the diameter thresholds, as shown in the figure, leads to larger cluster sizes progressing from left to right. This allows for estimating the upper bound of effective resistance between any two nodes based on the level where they share the same cluster index. For instance, nodes 5 and 9, with different cluster indices at levels (a), (b), and (c), share the same cluster index at level (d). Their resistance distance is bounded by the resistance diameter of cluster 2 at level (d).
@@ III-C1 Spectral Distortion Estimation
It has been demonstrated that inserting a new edge with a large spectral distortion  [^1] The spectral distortion of an edge is defined as the product of its edge weight and the effective resistance between its two end nodes [7, 9]. to the sparse graph will significantly increase its first few Laplacian eigenvalues [9, 7].
@@
The equation (6) proves that a higher spectral distortion $\Delta_{K}$ can be achieved by including edges with higher effective resistance $R(p,q)$ and higher edge weight $w_{p,q}$. Our algorithm efficiently estimates the effective resistance of the newly introduced edges, leveraging the node embedding information obtained from the setup phase. This enables sorting the newly introduced edges according to their estimated spectral distortion that leads to including the spectrally-critical edges to the sparse graph.
@@ III-C2 Spectral Similarity Estimation
The inGRASS algorithm applies an edge filtering process to exclude a newly added edge if there is already an existing edge in the sparse graph with a similar spectral distortion. This filtering process minimizes the introduction of additional edges while still influencing the spectral distortion. The algorithm selects a filtering level ($\mathcal{L}$) from the levels computed during the LRD decomposition phase with respect to a target condition number. For a target condition number $\kappa(L_{G},L_{H})=C$, the filtering level ($\mathcal{L}$) is chosen with the maximum number of nodes in a cluster equal to $\frac{C}{2}$. This threshold bounds the maximum spectral distortion among nodes within clusters (equivalent to the shortest path among the nodes within each cluster) to achieve the desired condition number.
Figure 3: (a) The original graph featuring three newly introduced edges highlighted in red. (b) The edge included into the sparse graph, marked in red, alongside the edges with adjusted weights, denoted in blue.
For a newly introduced edge, inGRASS only includes the edge if there is no existing edge in the sparse graph connecting the clusters to which those two nodes belong in the specified filtering level ($\mathcal{L}$). In the event of an existing edge connecting the clusters of those nodes in the filtering level ($\mathcal{L}$), the newly introduced edge is discarded, and the weight of the existing edge is adjusted by adding the weight of the newly added edge to it. Moreover, if the newly introduced edge connects two nodes within the same cluster in the filtering level ($\mathcal{L}$), the edge is discarded and its weight is proportionally distributed among the edges connecting the nodes within that cluster.
@@ III-D The Algorithm Flow and Complexity Analysis
The complete flow of the inGRASS algorithm is described in Algorithm 1 to efficiently update the sparse graph $H^{(0)}=(V,E^{\prime},w^{\prime})$ with $|V|=N$ after adding a set of weighted edges to the original graph $G^{(0)}$ to reach the target condition number $\kappa(L_{G},L_{H})=C$.
@@
Output: The updated sparse graph $H$.
@@ IV Experimental results
Table I provides a runtime comparison between GRASS and inGRASS setup times (measured in seconds) for various test cases. The runtime for the setup phase of inGRASS is mainly due to the LRD decomposition step for computing resistance embeddings for the initial sparse graph $H^{(0)}$. Note that this step is a one-time task, which can be leveraged for many subsequent update iterations involving streaming edge modifications. We observe that the inGRASS setup time is even faster than the GRASS time for most test cases as shown in Table I, where $|V|$ ($|E|$) denotes the number of nodes (edges) in the original graph $G^{(0)}$, respectively.
@@
Table II presents a thorough comparison of our incremental graph spectral sparsification algorithm, inGRASS, with GRASS and Random methods through $10$-iterative updates. The density ($D$), defined as $D:=\frac{|E|}{|V|}$, shows the density of the initial sparse graph $H^{(0)}$ and its density when all newly introduced edges are included. This emphasizes the importance of incremental spectral sparsification to prevent a substantial increase in sparse graph density. We maintain an initial density of $D=10\%$ for consistency. The relative condition number $\kappa(L_{G},L_{H})$ measures the spectral similarity between the original graph $G$ and the updated sparse graph $H$. A smaller relative condition number indicates higher spectral similarity. The column related to $\kappa(L_{G},L_{H})$ indicates how the condition number between $G^{(0)}$ and $H^{(0)}$ is perturbed when newly added edges are excluded from $H^{(0)}$, providing insights into edge generation. Columns ”GRASS-D,” ”inGRASS-D,” and ”Random-D” compare sparse graph density for a target condition number using each method. The target condition number is set to match the initial condition number between $G^{(0)}$ and $H^{(0)}$. For instance, the target condition number for ”G3 circuit” is set to $\kappa(L_{G^{(0)}},L_{H^{(0)}})=88$. Notably, inGRASS achieves the target condition number comparably to GRASS and significantly outperforms Random, accompanied by a runtime speedup exceeding $\frac{GRASS-T}{inGRASS-T}=200\times$ across all test cases through $10$-iterative updates.
@@ IV-B Robustness of the inGRASS Algorithm
The performance of our incremental graph spectral sparsification algorithm, inGRASS, is studied for the ”G2 circuit” dataset over different initial sparse graph densities. In Table III, the labeled column ”Density (D)” shows an initial density range of $6.5\%$ to $12.7\%$, with the sparse graph density set to $32\%$ when all edges are included. The $\kappa(L_{G},L_{H})$ column indicates the perturbation in the condition number between the original graph $G^{(0)}$ and the initial sparse graph $H^{(0)}$ when no newly added edges are included. The target condition number is set to the initial value (e.g. 56 for $D=12.7\%$). Both ”GRASS-D” and ”inGRASS-D” denote the sparse graph density for the same target condition number, with inGRASS showing comparable results to GRASS across various initial densities.
TABLE III: GRASS vs inGRASS densities across different initial densities in the sparse graph (“G2 circuit” test case).

### v2

@@ I Introduction
An implementation of our algorithm and the code for reproducing our experimental results are available online at https://github.com/Feng-Research/inGRASS.
@@ II-B Spectral Graph Decomposition
Furthermore, a similar spectral graph sparsification method is proposed in [9], which decomposes the graph into multiple sets of disjoint cycles using a multilevel spectral graph coarsening framework. By constructing a hierarchy of contracted graphs, the method identifies and adds spectrally-critical edges to the initial graph sparsifier.
@@ II-C Effective Resistance
$R_{eff}(p,q)=\sum\limits_{i=2}^{N}\frac{(u_{i}^{\top}b_{pq})^{2}}{\lambda_{i}},$ (2)
@@ III-A Overview of inGRASS Algorithm
This is a one-time operation that equips the inGRASS algorithm to iteratively update the initial graph sparsifier $H$ upon receiving streams of newly introduced edges in different iterations. In this phase, each node in the initial graph sparsifier $H^{(0)}$ is assigned a $\log N$-dimensional embedding vector by leveraging low-resistance-diameter (LRD) decomposition. A sparse data structure is then created that provides an efficient estimate of the resistance distance among different pair of nodes in the initial graph sparsifier $H^{(0)}$.
@@ III-B1 Scalable Estimation of Effective Resistances
Figure 2: A 4-level resistance embedding of the initial graph sparsifier achieved through the proposed LRD decomposition. Since the embedding vectors for nodes $5$ and $9$ are $[3,3,3,2]^{\top}$ and $[6,2,2,2]^{\top}$, the effective-resistance distance between them is bounded by the resistance diameter of cluster $2$ shown in (d).
@@
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L_{G}\tilde{u}_{i}},$ (3)
@@ III-B2 Multilevel Resistance Embedding
We utilize a multilevel low-resistance-diameter (LRD) decomposition framework for embedding nodes of the initial graph sparsifier $H^{(0)}$ into an $O(\log N)$-dimensional space [17]. The embedding process iteratively follows these steps: (S1) efficiently estimate the effective resistance of edges using (3); (S2) contract edges starting from low-resistance-diameter clusters (cluster diameters are initialized to 0 for all nodes in the first level); (S3) replace a contracted edge with a supernode and adjust the resistance diameter of the clusters accordingly.
@@ III-B3 Multilevel Sparse Data Structure
Leveraging the $O(\log N)$-dimensional node embedding vectors obtained in the previous step, we construct a specialized sparse data structure for efficient access to the node clustering indices at different levels. This facilitates efficient storage and retrieval of cluster connectivity information, enabling the identification of spectrally-critical and spectrally-unique edges among the newly added edges. The sparse data structure is promptly updated upon the addition of a newly introduced edge to the graph sparsifier.
Fig. 2 illustrates the computation of node embedding vectors through multilevel LRD decomposition of the initial graph sparsifier. Increasing the diameter thresholds, as shown in the figure, leads to larger cluster sizes progressing from left to right. This allows for estimating the upper bound of effective resistance between any two nodes based on the level where they share the same cluster index. For instance, nodes 5 and 9, with different cluster indices at levels (a), (b), and (c), share the same cluster index at level (d). Their resistance distance is bounded by the resistance diameter of cluster 2 at level (d).
@@ III-C1 Spectral Distortion Estimation
It has been demonstrated that inserting a new edge with a large spectral distortion  [^1] The spectral distortion of an edge is defined as the product of its edge weight and the effective resistance between its two end nodes [7, 9]. to graph sparsifier will significantly increase its first few Laplacian eigenvalues [9, 7].
@@
The equation (6) proves that a higher spectral distortion $\Delta_{K}$ can be achieved by including edges with higher effective resistance $R(p,q)$ and higher edge weight $w_{p,q}$. Our algorithm efficiently estimates the effective resistance of the newly introduced edges, leveraging the node embedding information obtained from the setup phase. This enables sorting the newly introduced edges according to their estimated spectral distortion that leads to including the spectrally-critical edges to the graph sparsifier.
@@ III-C2 Spectral Similarity Estimation
The inGRASS algorithm applies an edge filtering process to exclude a newly added edge if there is already an existing edge in the graph sparsifier with a similar spectral distortion. This filtering process minimizes the introduction of additional edges while still influencing the spectral distortion. The algorithm selects a filtering level ($\mathcal{L}$) from the levels computed during the LRD decomposition phase with respect to a target condition number. For a target condition number $\kappa(L_{G},L_{H})=C$, the filtering level ($\mathcal{L}$) is chosen with the maximum number of nodes in a cluster equal to $\frac{C}{2}$. This threshold bounds the maximum spectral distortion among nodes within clusters (equivalent to the shortest path among the nodes within each cluster) to achieve the desired condition number.
Figure 3: (a) The original graph featuring three newly introduced edges highlighted in red. (b) The edge included into the graph sparsifier, marked in red, alongside the edges with adjusted weights, denoted in blue.
For a newly introduced edge, inGRASS only includes the edge if there is no existing edge in the graph sparsifier connecting the clusters to which those two nodes belong in the specified filtering level ($\mathcal{L}$). In the event of an existing edge connecting the clusters of those nodes in the filtering level ($\mathcal{L}$), the newly introduced edge is discarded, and the weight of the existing edge is adjusted by adding the weight of the newly added edge to it. Moreover, if the newly introduced edge connects two nodes within the same cluster in the filtering level ($\mathcal{L}$), the edge is discarded and its weight is proportionally distributed among the edges connecting the nodes within that cluster.
@@ III-D The Algorithm Flow and Complexity Analysis
The complete flow of the inGRASS algorithm is described in Algorithm 1 to efficiently update the graph sparsifier $H^{(0)}=(V,E^{\prime},w^{\prime})$ with $|V|=N$ after adding a set of weighted edges to the original graph $G^{(0)}$ to reach the target condition number $\kappa(L_{G},L_{H})=C$.
@@
Output: The updated graph sparsifier $H$.
@@ IV Experimental results
Table I provides a runtime comparison between GRASS and inGRASS setup times (measured in seconds) for various test cases. The runtime for the setup phase of inGRASS is mainly due to the LRD decomposition step for computing resistance embeddings for the initial graph sparsifier $H^{(0)}$. Note that this step is a one-time task, which can be leveraged for many subsequent update iterations involving streaming edge modifications. We observe that the inGRASS setup time is even faster than the GRASS time for most test cases as shown in Table I, where $|V|$ ($|E|$) denotes the number of nodes (edges) in the original graph $G^{(0)}$, respectively.
@@
Table II presents a thorough comparison of our incremental graph spectral sparsification algorithm, inGRASS, with GRASS and Random methods through $10$-iterative updates. The density ($D$), defined as $D:=\frac{|E|}{|V|}$, shows the density of the initial graph sparsifier $H^{(0)}$ and its density when all newly introduced edges are included. This emphasizes the importance of incremental spectral sparsification to prevent a substantial increase in graph sparsifier density. We maintain an initial density of $D=10\%$ for consistency. The relative condition number $\kappa(L_{G},L_{H})$ measures the spectral similarity between the original graph $G$ and the updated graph sparsifier $H$. A smaller relative condition number indicates higher spectral similarity. The column related to $\kappa(L_{G},L_{H})$ indicates how the condition number between $G^{(0)}$ and $H^{(0)}$ is perturbed when newly added edges are excluded from $H^{(0)}$, providing insights into edge generation. Columns ”GRASS-D,” ”inGRASS-D,” and ”Random-D” compare graph sparsifier density for a target condition number using each method. The target condition number is set to match the initial condition number between $G^{(0)}$ and $H^{(0)}$. For instance, the target condition number for ”G3 circuit” is set to $\kappa(L_{G^{(0)}},L_{H^{(0)}})=88$. Notably, inGRASS achieves the target condition number comparably to GRASS and significantly outperforms Random, accompanied by a runtime speedup exceeding $\frac{GRASS-T}{inGRASS-T}=200\times$ across all test cases through $10$-iterative updates.
@@ IV-B Robustness of the inGRASS Algorithm
The performance of our incremental graph spectral sparsification algorithm, inGRASS, is studied for the ”G2 circuit” dataset over different initial graph sparsifier densities. In Table III, the labeled column ”Density (D)” shows an initial density range of $6.5\%$ to $12.7\%$, with the graph sparsifier density set to $32\%$ when all edges are included. The $\kappa(L_{G},L_{H})$ column indicates the perturbation in the condition number between the original graph $G^{(0)}$ and the initial graph sparsifier $H^{(0)}$ when no newly added edges are included. The target condition number is set to the initial value (e.g. 56 for $D=12.7\%$). Both ”GRASS-D” and ”inGRASS-D” denote the graph sparsifier density for the same target condition number, with inGRASS showing comparable results to GRASS across various initial densities.
TABLE III: GRASS vs inGRASS densities across different initial densities in the graph sparsifier (“G2 circuit” test case).
@@ V Conclusion
Acknowledgment
This work is supported in part by the National Science Foundation under Grants CCF-2205572, CCF-2021309, CCF-2212370, and CCF-2011412.